OpenStudy (anonymous):

Compare and contrast Sin and Cos functions. Hint: Period, shape, minimum point, maximum point

OpenStudy (anonymous):

Do you mean \(\mbox{sin x}\) and \(\mbox{cos x}\)?

OpenStudy (anonymous):

@tinybookworm the functions so I guess, yeah

OpenStudy (anonymous):

The period of both \(\mbox{sin x}\) and \(\mbox{cos x}\) is \(2\pi\).

OpenStudy (anonymous):

I have that.

OpenStudy (anonymous):

Next, you may already know how they look like, the sinusoid

OpenStudy (anonymous):

@marshallinwashington can you help?

OpenStudy (anonymous):

@tinybookworm What is the sinusoid again?

OpenStudy (anonymous):

It is the sine wave form, like this

OpenStudy (anonymous):

Thanks. Anymore?

OpenStudy (anonymous):

And the max and min points: In the function \(\mbox{y = sin x}\), the max points are \(\large (\frac{\pi}{2},1)\), \(\large (\frac{5\pi}{2},1)\), \(\large (\frac{9\pi}{2},1)\), \(\large (\frac{13\pi}{2},1)\), and so on. The min points are \(\large (\frac{3\pi}{2},-1)\), \(\large (\frac{7\pi}{2},-1)\), \(\large (\frac{11\pi}{2},-1)\), \(\large (\frac{15\pi}{2},-1)\), and so on. In the function \(\mbox{y = cos x}\), the max points are \((0,1)\), \(\large (2\pi,1)\), \(\large (4\pi,1)\), \(\large (6\pi,1)\), and so on. The min points are \((\pi,-1)\), \(\large (3\pi,-1)\), \(\large (5\pi,-1)\), \(\large (7\pi,-1)\), and so on.

OpenStudy (anonymous):

As you can see in the picture, the wave is repeated so there are infinite max points and min points

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

You are welcome